### v1

# Sorting inversion sequences

Abstract
We consider the avoidance of patterns in inversion sequences that relate sorting via sorting machines including data structures such as pop stacks and stacks. Such machines have been studied under a variety of additional constraints and generalizations, some of which we apply here. We give the classification of several classes of sortable inversion sequences in terms of pattern avoidance. We are able to provide an exact enumeration of some of the sortable classes in question using both classical approaches and a more recent strategy utilizing generating trees.
1.   Introduction
An inversion in a permutation $\pi=\pi_{1}\pi_{2}\pi_{3}\cdots\pi_{n}$ is a pair of entries $\pi_{i},\pi_{j}$ where $i<j$, but $\pi_{i}>\pi_{j}$. An inversion sequence of length $n$ is a word $e=e_{0}\dotsb e_{n}$ which satisfies for each $i\in[n]=\{0,1,\ldots,n\}$ the inequalities $0\leq e_{i}\leq i$. These sequences correspond to permutations via a variant of the Lehmer code [17, 18] by mapping each entry $\pi_{j}$ of the permutation $\pi$ to $e_{j}$ where $e_{j}$ is the number of inversions $\pi_{j}$ is a part in $\pi$ as the second, smaller entry of the pair. The set of inversion sequences of length $n$ is denoted ${\mathbf{I}}_{n}$.
The study of pattern avoidance in inversion sequences began with work by Corteel, Martinez, Savage, and Weselcouch [8] and Shattuck and the first author [21]. Since then, extensive work has been done in this area, including recent work by Kotsireas, Yıldırım, and the first author [16] that introduces an algorithmic technique involving generating trees to determine enumeration of many pattern classes of inversion sequences (for example, see [6, 19, 20, 30] and the references therein). We utilize this algorithm ourselves in Section 4 to obtain a generating function for the number of inversion sequences sortable by two particular machines.
However, there remain many open problems when it comes to characterizing and enumerating the inversion sequences that avoid a class of consisting of a small number of relatively short (length less than five) patterns. Our interest is in the characterization and enumeration of inversion sequences sortable by stacks and pop stacks. We also study the characterization of inversion sequences sortable by data structures obtained by generalizing or further restricting the pop stacks.
A stack is a last-in, first-out sorting device with push and pop operations. Knuth [15] showed the permutation $\pi$ can be sorted (that is, by applying push and pop operations to the sequence $\pi_{1},\dots,\pi_{n}$ one can output the identity $1,\dots,n$) exactly if $\pi$ avoids the permutation $231$.
Definition 1.1.
A permutation $\pi=\pi_{1}\pi_{2}\dots\pi_{n}\in S_{n}$ is said to contain a permutation $\sigma=\sigma_{1}\sigma_{2}\ldots\sigma_{k}$ if there exist indices $1\leq\alpha_{1}<\alpha_{2}<\ldots<\alpha_{k}\leq n$ such that $\pi_{\alpha_{i}}<\pi_{\alpha_{j}}$ if and only if $\sigma_{i}<\sigma_{j}$. Otherwise, we say $\pi$ avoids $\sigma$.
These definitions of containment and avoidance have been naturally extended to words (and specifically inversion sequences) with the added allowance of entries being equal to one another.
Example 1.2.
The permutation $\pi=241563$ contains $312$ since the $4,1,3$ appear in the same relative order as $3,1,2$. However, $\pi$ avoids $321$ since there is no decreasing subsequence of length three in $\pi$.
Similarly, the inversion sequence $\tau=0021104$ contains $1002$ from $2,1,1,4$, but avoids $201$.
Describing the permutations that are sortable by these machines is a point of interest in the field. To this end, it often possible to describe the sortable machine in terms of its basis.
Definition 1.3.
A permutation class is a downset of permutations under the containment order. Every permutation class can be specified by the set of minimal permutations which are not in the class called its basis. For a set $B$ of permutations, we denote by $\operatorname{Av}(B)$ the class of permutations which do not contain any element of $B$.
As mentioned above, the stack-sortable permutations are the class $\operatorname{Av}(231)$.
A restricted version of a stack that has received a fair amount of attention is a pop stack, first introduced by Avis and Newborn [4]. A pop stack has the same push and pop operations as a stack, but whenever a pop operation occurs, every entry in the pop stack is popped immediately. One foundational result is that the permutations that can be sorted by a single pop stack are exactly those that avoid $231$ and $312$. Some later works involving pop stacks include [1, 3, 7, 9, 12, 14, 25, 28].
Atkinson [2] explored a variation of a stack called an $(r,s)$-stack which allowed for an extension of the traditional push and pop operations. Specifically, the first index of an $(r,s)$-stack allowed for one to push an entry not only to the top of the stack, but to any of the top $r$ spots. Similarly, the second index expanded the options of the pop operation to allow one to pop any of the top $s$ entries from the stack. Thus a $(1,1)$-stack is simply a traditional stack. Atkinson was able to characterize these permutations when one of $r,s$ was fixed at $1$ and when $r=s=2$. Natural variations on stacks arise when studying inversion sequences that are similar in flavor to these generalized stacks.
Murray [10] considered restricting the depth of a stack, that is the maximum number of entries allowed in a stack at any given time, in the context of generating permutations with two stacks in series. Murray later continued this study with Lee and Rechnitzer [13] and then extended to sorting and pop stack sorting with Goh [11, 12]. While this work focused on a permutation sorting (or generating) machine consisting of two stacks in series, we will consider a similar restriction for a single stack or pop stack when sorting inversion sequences.
2.   Stack sorting inversion sequences
We give the standard optimal stack sorting algorithm in Algorithm 1.
Algorithm 1 Stack Sorting Algorithm
$p=p_{1}p_{2}\cdots p_{n}$ is a permutation of length $[n]$.
$S$ is a stack.
$O$ is an empty array.
$i=1$
procedure StackSort($p,n$)
     while $i<n+1$ do
         if $S$ is empty then
              push $p_{i}$ to the top of $S$
              $i=i+1$
         else if $p_{i}<$ top element of $S$ then
              push $p_{i}$ to the top of $S$
              $i=i+1$
         else
              pop the top entry of $S$ to the end of $O$
         end if
     end while
return $O$
end procedure
For a word $w$ to be stack sortable, $w$ must avoid $120$, the same as avoiding $231$ to be able to sort permutation. The repetition allowed in words does not change this requirement from being both necessary and sufficient and so the proof does not change either. It is included below for completeness.
Proposition 2.1.
The words that are stack sortable are exactly the words that avoid $120$.
Proof.
The first $2$ of a $120$ pattern forces the smaller entry $1$ to the output before the 00 even enters the stack which creates an inversion with the $1,0$. Hence words containing a $120$ pattern are not stack sortable.
Next consider a word $w$ that is not stack sortable. The output $S(w)$ must contain an inversion, say an entry $b>a$ that appears before $a$. For $a$ to not have exited the stack before $b$, we know $a$ appears after $b$ in $w$. Further, for $b$ to have been forced out of the stack before $a$ entered, there must have been a larger entry $c$ that appeared after $b$, but before $a$ to force $b$ out before $a$ could enter the stack. Thus $w$ contains a $120$ pattern $bca$. ∎
As inversion sequences are words with added restrictions, Proposition 2.1 also gives us a basis for the stack sortable inversion sequences.
Corollary 2.2.
The inversion sequences that are stack sortable are exactly those that avoid $120$.
The enumeration of these sortable inversion sequences has been studied, but thus far remains open. See [21, Section 4] and [8] for initial study and further work on avoiding $120$ in [30].
3.   Pop stack sorting inversion sequences
The optimal pop stack sorting algorithm is mostly the same procedure as that for stacks, but when a pop stack is popped, all entries must exit the stack as described in Algorithm 2.
Algorithm 2 Pop Stack Sorting Algorithm
$p=p_{1}p_{2}\cdots p_{n}$ is a permutation of length $[n]$.
$PS$ is a pop stack.
$O$ is an empty array.
$i=1$
procedure PopStackSort($p,n$)
     while $i<n+1$ do
         if $PS$ is empty then
              push $p_{i}$ to the top of $PS$
              $i=i+1$
         else if $p_{i}<$ top element of $PS$ then
              push $p_{i}$ to the top of $PS$
              $i=i+1$
         else
              while $PS$ is nonempty do
                  pop the top entry of $PS$ to the end of $O$
              end while
         end if
     end while
return $O$
end procedure
Note that $120,201$ are the patterns whose avoidance completely determines the pop stack sortability of a permutation. However, due to the repetition allowed in words, a copy of a larger element already in the stack can force the entire stack to be popped while there is still a copy of a smaller entry appearing later in the permutation.
Theorem 3.1.
The words that are pop stack sortable are exactly the words that avoid $120,201,$ and $1010$.
Proof.
One can check that each of $120,201,$ and $1010$ are not pop stack sortable. As such, a word that contains any of these patterns cannot be sorted by a pop stack.
Conversely, suppose $w$ is not pop stack sortable. Consider what happens to force an entry to the output that is larger than an entry that appears later in the permutation. Using Algorithm 2, we had an entry $d$ that appears (immediately) after a smaller entry $b$ which forces $b$ out before $d$ can enter. If $b$ is the larger entry forced out, then there exists a smaller entry $a$ after $d$ and $bda$ form a $120$ pattern. If not, then there exists an entry $c>b$ that entered the pop stack before $b$ and there is an entry $a$ that appears after $d$ such that $c>a\geq b$. Note if $a>b$, then $cba$ form a $201$ pattern. Otherwise, if $a=b$, possible patterns $cbda$ are $2010,10$ 2010 , 1020, but only $1010$ does not contain one of the smaller forbidden patterns $120,201$. ∎
As inversion sequences are words with added restrictions, Theorem 3.1 also gives us a basis for the pop stack sortable inversion sequences.
Corollary 3.2.
The inversion sequences that are pop stack sortable are exactly those that avoid $120,201,$ and $1010$.
The pop stack sortable permutations are also known as layered permutations which are permutations of the form $\pi=\sigma_{1}\sigma_{2}\cdots\sigma_{k}$ where each layer $\sigma_{i}$ is a maximal length contiguous decreasing sequence such that all of the entries of $\sigma_{i}$ are less than all of the entries of $\sigma_{i+1}$. For example, the permutation $\pi=543216987$ is a layered permutation with layers $\sigma_{1}=54321,\sigma_{2}=6,\sigma_{3}=987$ as shown in Figure 1.
Figure 1: The layered permutation $\pi=543216987$
We can define layered words similarly as follows.
Definition 3.3.
A layered word $w$ is such that $w=\tau_{1}\tau_{2}\cdots\tau_{k}$ where each $\tau_{i}$ is
  1. 1.
    maximal in length,
  2. 2.
    weakly decreasing, and
  3. 3.
    such that the last (weakly smallest) entry of $\tau_{i+1}$ is at least as large as the first (weakly largest) entry of $\tau_{i}$ for all $i$.
For example, the word $w=211033275$ is a layered word with layers $2110,332,75$ shown in Figure 2.
Figure 2: The layered word $\pi=211033275$
It can then be seen that these layered words are exactly the pop stack sortable words.
Proposition 3.4.
Layered words are exactly the words that avoid $120,201,1010$.
Proof.
Suppose $w$ is layered. Then $w$ cannot contain $120$ because if not, the layer containing the 00 of the pattern must appear after the layer containing the $1$ which violates the third condition. The same argument can be made for the $1$ of the $201$ pattern appearing in a later consecutive decreasing pattern than that of the $2$ and once again for the second 00 of the $1010$ pattern appearing in a later consecutive decreasing pattern than that of the first $1$.
Now suppose $w$ avoids $120,201,1010$. Then $w$ is pop stack sortable. Consider the output from sorting $w$ broken into segments based on the pops. Reversing those segments reverts the word back to $w$. Further, by the algorithm, these breaks occur at exactly the strict ascents in $w$. The segments are then weakly decreasing and such that the smallest entry of any segment is at least as large as the largest entry of the previous segment for $w$ to be sortable and also maximal in length. Hence $w$ is layered. ∎
When restricting to inversion sequences, we are able to describe the generating tree which we describe in subsection 4.2. There are also some partial results we can obtain combinatorially based on the number of layers (which is also the number of pops needed to sort the sequence). Recall the Eulerian numbers $E(n,k)$ count the number of permutations of length $n$ with $k-1$ descents.
Proposition 3.5.
The number of pop stack sortable inversion sequences of length $n$ with $k=1,2$ layers are counted by $E(n,k)$.
Proof.
In fact, the pop sortable inversion sequences of length $n$ with $k=1,2$ layers correspond exactly to the permutations of length $n$ with $k$ descents by the variant of Lehmer code used to create inversion sequences. In the trivial case of $k=1$, a permutation of length $n$ with $k-1=0$ descents is the identity permutation which is the only permutation of length $n$ to have an inversion sequence with one layer, namely the inversion sequence consisting of all 00s.
Now consider the case when $k=2$. A permutation $\pi$ with exactly one descent, say at index $i$, will have an inversion sequence that begins with $i$ 0s which will be the first layer. Since the permutation $\pi$ is such that $\pi_{i+1}$ is the smaller entry in an inversion with at least $\pi_{i}$, the corresponding inversion sequence now has a positive entry in position $i+1$. As the remainder of the entries of $\pi$ are all in increasing order, each subsequent entry of $\pi$ can be the smaller entry of at most as many inversions as the previous entry. Thus, the rest of the inversion sequence is in weakly decreasing order meaning that the entries from position $i+1$ to position $n$ form the second and final layer. ∎
As not all inversion sequences are pop stack sortable, this correspondence can clearly not continue indefinitely. Indeed, a permutation with even $k-1=2$ descents is not guaranteed a corresponding pop stack sortable inversion sequence. For example, the permutation $\pi=3214$ has the inversion sequence $e(\pi)=0120$ which is not sortable by a pop stack.
At the other extreme, only one inversion sequence of any given length $n\geq 1$ can have $n$ layers, namely the increasing sequence $012\ldots(n-1)$. Reducing the number of layers by one to $n-1$ layers gives a set of inversion sequences enumerated by the Tetrahedral numbers, sequence number A000292 in OEIS [27].
Proposition 3.6.
The number of pop stack sortable inversion sequences of length $n\geq 1$ with $n-1$ layers is $\binom{n+1}{3}$.
Proof.
A pop stack sortable inversion sequence of length $n\geq 1$ with $n-1$ layers will have exactly one entry $e_{i}$ as the second entry of a layer where $i\in\{2,3,4,\ldots,n\}$. It must be the case that $e_{m}=m-1$ for all $m<i$ for those entries to have all started new layers. This means $e_{i}=e_{i-1}=i-2$ or $e_{i}=e_{i-2}=i-3$. Further, we know $e_{i},e_{i+1},e_{i+2},\ldots,e_{n}$ are in increasing order as they are all in different layers.
In the case that $e_{i}=i-2$, since we must have $i-1\leq e_{i+1}<e_{i+2}<\cdots<e_{n}\leq n-1$, we can select up to one of these later entries to be two greater than the previous entry. Consider positions $\{2,3,4,\ldots,n,n+1\}$ where we pick two positions; the first to be the position of the second entry in the layer of length two and the second to be the position of the entry that is two more (as opposed to one more) than the previous. In the case that $n+1$ is one of the selected options, then we do not have any entry that is two more than the one before it. Hence there are $\binom{n}{2}$ such inversion sequences.
Otherwise, $e_{i}=i-3$ and so $i\geq 3$. Now because $i-2\leq e_{i+1}<e_{i+2}<\cdots<e_{n}\leq n-1$, up to two of these later entries could be two greater than the previous entry or one of these later entries could be three greater than the previous entry. For the cases when no entry is three more than the previous entry, we have $\binom{n-2}{1}+\binom{n-2}{2}+\binom{n-2}{3}$ ways to select the second entry of the length two layer and the entries that are two greater than the entry they follow. In the case where we have one entry that is three more than the previous entry, there are $\binom{n-2}{2}$ ways to select this entry and the entry that was the second in the longest layer. This gives us a total of $\binom{n-2}{1}+2\binom{n-2}{2}+\binom{n-2}{3}=\binom{n}{3}$ via a generalization of Pascal’s Identity.
By another application of Pascal’s Identity, there are $\binom{n}{2}+\binom{n}{3}=\binom{n+1}{3}$ pop stack sortable inversion sequences of length $n\geq 1$ with $n-1$ layers. ∎
One last special case is when the sortable inversion sequence contains only 00s and $1$s and thus only needs to avoid $1010$, thereby having at most three layers. The number of such sequences is the number of compositions of $n$ into at most four parts because we must begin with a 00 and can only alternate values at most three times after that. These numbers (with an offset of one) are better known as the “Cake numbers" [29], sequence number A000125 in OEIS [27], which count the maximum number of slices of a cylindrical cake one can have after making $n-1$ planar cuts perpendicular to the top circular surface.
Proposition 3.7.
The number of pop sortable inversion sequences of length $n$ only containing $0,1$ is $(n-1)$st Cake number: $\binom{n-1}{0}+\binom{n-1}{1}+\binom{n-1}{2}+\binom{n-1}{3}$.
3.1 Restricting pop stack depth
When extending the notion of the depth of a stack to words, we could continue to use the definition in terms of the number of elements allowed in the stack as introduced by Murray [10]. However, given the repetition that distinguishes words from permutations, we will instead use the following definition that is also consistent with the original definition applied to permutations.
Definition 3.8.
The depth of a stack (or pop stack) will be the number of entries with distinct values allowed in the stack (or pop stack) at any stage.
For example, a stack of depth three under this definition would still only allow three entries of a permutation to be in the stack at one time. However, this same restricted stack could allow arbitrarily many entries of a word in the stack, provided that no more than three distinct letters were represented at any time. We begin by considering pop stacks of limited depth.
Proposition 3.9.
A word is sortable by a pop stack of depth $k$ if and only if avoids $120,201,1010$ and also $k(k-1)\cdots 10$.
Proof.
As before any word $w$ must avoid $120,201,1010$ to be pop stack sortable. Additionally, a decreasing sequence that is too long to fit in a limited depth pop stack is also not sortable.
Conversely, if $w$ is not pop stack sortable by a pop stack of depth $k$ and avoids $120,201,1010$, then the problem is a result of the limitation of the pop stack’s capacity. Hence $w$ must contain a $k(k-1)\cdots 10$ pattern. ∎
Corollary 3.10.
An inversion sequence is sortable by a pop stack of depth $k$ if and only if avoids $120,201,1010$ and also $k(k-1)\cdots 10$.
In the case of depth one, the number of sortable inversion sequences is the number of weakly increasing inversion sequences which are known to be enumerated by the Catalan numbers due to Martinez and Savage [24]. The proof shows these inversion sequences to correspond to the $213$ avoiding permutations enumerated by Simion and Schmidt [26].
Theorem 3.11.
(Martinez and Savage) The number of inversion sequences of length $n$ which avoid $10$ and thus are sortable by a pop stack of depth $1$ is the $n$th Catalan number $C_{n}=\frac{\binom{2n}{n}}{n+1}$.
For a pop stack of depth two, the sortable inversion sequences would need to avoid $120,201,210,1010$. The first few terms of the enumeration sequence are: $1,1,2,6,23,100,471,2349$. We can define the generating tree for these inversion sequences (using the algorithm introduced in [16]) as having a root $a_{1}$ and satisfying the following rules
$a_{m}$ $\rightsquigarrow a_{m+1},b_{m,1},\ldots,b_{m,m},\quad m\geq 1,$
$b_{m,j}$ $\rightsquigarrow(c_{m+1-j})^{j},b_{m+1,j},b_{m+1-j,1},\ldots,b_{m+1-j,m+1-j},\quad 1\leq j\leq m,$
$c_{m}$ $\rightsquigarrow c_{m+1},a_{m+2},b_{m+1,1},\ldots,b_{m+1,m+1},\quad m\geq 1,$
where $a_{m}=0^{m}$, $b_{m,j}=0^{m}j$ and $c_{m}=0^{m}10$. However, it is hard to get an explicit formula by translating this generating tree to system of equations and is even more difficult for larger values of $k$.
Another approach is to count the inversion sequences recursively directly, but this too is not as nice as we would hope. One recursive characterization is given below.
A weakly decreasing word of length $n$ has depth $k$ if it has the form $\{a_{1},...a_{1},a_{2},...a_{2},...,a_{k+1},...a_{k+1}\}$ where for each $i$, $a_{i}>a_{i+1}\geq 0$ (so there are $k$ descents). Let $WD_{k}(n,a)$ be the set of all such words whose first value is $a$ and whose depth is at most $k$.
Lemma 3.12.
For all $n,k\in\mathbb{Z}^{+}$, with $k<n$
$|WD_{k}(n,a)|=1+\sum_{j=1}^{k}\binom{n-1}{j}\binom{a-1}{j-1}$ (1)
Note we assume $\binom{n}{k}=0$ if $k>n$. The result follows from choosing $j$ locations for a descent, then $j-1$ values for the weakly decreasing sequence, and one for the sequence with no descents.
We say a word is a layered weakly decreasing word of depth at most $k$ and length $n$ if it can be written as a sequence of subwords $\{w_{1},w_{2},...\}$ where each $w_{i}\in WD_{k}(m_{i},a_{i})$ with $\sum_{i}m_{i}=n$, and for each $i$, $a_{i}\geq 0$, and $min(w_{i+1})\geq max(w_{i})$ (i.e. the largest (first) value of each word is bounded by the smallest value of the succeeding word). Further, we consider the words of this type which are also inversion sequences, i.e. $max(w_{i})\leq\sum_{j=1}^{i-1}m_{j}$. Let $LDI_{k}(n)$ be the set of all such layered depth at most $k$ inversion sequences of length $n$.
Lemma 3.13.
The set of sortable inversion sequences by a depth $k$ pop stack is exactly $LDI_{k}(n)$.
The proof is simply to note that when the stack is popped, there are at most $k$ distinct values, and they must be less than any later sequence values.
Given an element $\{w_{1},w_{2},...w_{t}\}\in LDI_{k}(n)$ as above with each value of $w_{i+1}$ bounded below by the largest value in $w_{i}$, subtracting $a_{i}$ from every value in $w_{i+1}$ yields an element of $WD_{k}(m_{i},a_{i+1}-a_{i})$.
As a result we get the following formula.
Proposition 3.14.
For all $n,k>0$,
$|LDI_{k}(n)|=\sum_{m_{1}+m_{2}+...m_{t}=n,m_{i}\geq 0}\sum_{a,1,a_{2},...a_{t}}^{*}|WD(m_{i},a_{i})|,$ (2)
where the $\sum^{*}$ means to only sum on values of ${0\leq a_{j}\leq\bigg{(}\sum_{h=1}^{j-1}m_{h}-a_{h}\bigg{)}-1}$.
Since the first word in an element of $LDI_{k}(n)$ must be a string of zeros, we let $LDI_{k}(n,m_{1})$ denote the elements of $LDI_{k}(n)$ that begin with $m_{1}$ zeros followed by a non-zero value. Further, let $LDI_{k}(n,m_{1},a_{2})$ is the subset of these layered decreasing inversion sequences which begin with $m_{1}$ zeros, and first non-zero value $a_{2}$ (and $a_{2}\leq m_{1}$). We now decompose our sequence of words as $\{\overline{0}_{m_{1}},w_{2},\hat{w}\}$ where $\hat{w}=\{w_{3},...w_{t}\}$. Now $w_{2}\in WD_{k}(m_{2},a_{2})$ for some $0\leq m_{2}\leq n-m_{1}$, $a_{2}>0$. If we subtract $a_{2}$ from all the elements in $\hat{w}$ the result is in $LDI(n-m_{2}-m_{1},m_{1}+m_{2}-a_{2},a)$ for some $a$ with $a\leq m_{1}+m_{2}-a_{2}$. Note that the number of starting zeros guarantees that adding $a_{2}$ to all of these entries will still yield an element of $LDI_{k}(n)$.
Hence we get the formula,
Proposition 3.15.
For all $n,k,m_{1}\in\mathbb{Z}^{+}$ with $m_{1}<n$,
$LDI_{k}(n,m_{1})=\sum_{m_{2}=0}^{n-m_{1}}\sum_{a_{2}\leq m_{1}}WD_{k}(m_{2},a_{2})\sum_{a\leq m_{1}+m_{2}-a_{2}}LDI_{k}(n-m_{1}-m_{2},m_{1}+m_{2}-a_{2},a).$ (3)
Both of the above formulas are rather difficult to use for computation. However they are presented to give some insight into directly enumerating these inversion sequences.
For completeness, we give some characterizations of words sortable by stacks of depth $k$.
Proposition 3.16.
A word is sortable by a stack of depth $k$ if and only if avoids $120$ and also $k(k-1)\cdots 10$.
Corollary 3.17.
The number of words of length $n$ on an alphabet $[k]$ sortable by a stack of depth $1$, i.e. avoid $10$ is $\binom{n+k-1}{k-1}$.
Proof.
These weakly increasing words can be thought of as weak compositions of $n$ into $k$ parts where each part represents the next largest letter. ∎
The depth 2 stack sortable words are enumerated by Burstein [5] in the context of avoiding $120,210$.
Theorem 3.18.
(Burstein) The number of words of length $n$ on an alphabet $[k]$ avoiding $120,210$ and thus sortable by a stack of depth $2$ is given by
$\frac{1-(-1)^{k}}{2}+2^{n}\sum_{i=0}^{\lfloor\frac{k-2}{2}\rfloor}\binom{n+k-3-2i}{n-1}.$
Finally, we note Kotsireas, Yıldırım, and the first author [16] give a functional equation for the inversion sequences of length $n$ avoiding $120,210$ and thereby sortable by a stack of depth $2$.
3.2 Generalizing pop stacks
We can also expand the ability of our sorting machine using a similarly defined version of Atkinson’s [2] extension of push and pop operations. Specifically, we will define a $(r,1)$-pop stack that is also modified to take into account the repetition of letters in inversion sequences (and more generally words). In Atkinson’s work, an $(r,1)$-stack would have allowed for push operations to push entries into the stack in any of the top $r$ positions. We can again naturally redefine generalized stacks for words in a way that is consistent with Atkinson’s work on permutations.
Definition 3.19.
Let an $(r,1)$-stack allow the push operation to push an entry to any of the top $r$ “value positions” in the stack where consecutive entries of the same value in the stack are grouped together. Note that the same definition could be extended to the broader notion of an $(r,s)$ stack where we allow the pop operation to be defined in terms of these same value positions.
Example 3.20.
[…]
Figure 3: A $(3,1)$-stack
Varying the restrictions of depth and generalizations of the push operation as described above, we get a family of stacks and a family of pop stacks. There are some variations that no longer give us new machines though. Specifically, if our stack or pop stack has limited depth, having the option to push entries to (or pop entries from) spaces that do not exist does add any sorting power.
Proposition 3.21.
For a $(r,1)$-stack or $(r,1)$-pop stack of depth $d$, if $r>d$, the machine sorts the same words as if $r=d$.
As such, we will restrict $r\leq d$.
We will first consider the case where $r=d=2$.
The $(2,1)$-pop stack of depth two has the nice characteristic where it sorts words (and permutations) that avoid three patterns of length three.
Theorem 3.22.
A $(2,1)$-pop stack of depth two sorts exactly the words that avoid $120,201,210$.
Proof.
Because the $(2,1)$-pop-stack we are using only has depth two, any pattern of length three with three distinct values will have at least the first entry popped out before the third entry can enter. For all of $120,201,210$, this means the larger first entry is popped out before the smaller last entry enters the stack. Thus no word containing $120,201,210$ is sortable by a $(2,1)$-pop stack of depth two.
Now consider a word $w$ that cannot be sorted by a $(2,1)$-pop stack of depth two. If a larger entry was forced out before a smaller entry could be pushed into the stack due to the limited depth, there are two ways this could have happened. One is that two distinct larger entries appeared before a smaller entry, that is $w$ contains a $120$ or $210$ pattern. Otherwise, we have a large entry that is below a small entry in the pop stack, and our third entry waiting to enter is between the small and large entry in value. Because of the $(2,1)$ generalization, in theory, the first two entries could appear in either order in $w$ to form either a $021$ or a $201$ pattern. However, in the case of a $021$ pattern, the 00 entry could be popped out before pushing the $2$ into the pop stack, so only the $201$ pattern presents a problem. ∎
Note: Algorithm 3 gives a procedure for sorting with a $(2,1)$-pop stack of depth $2$. Because there are only two possible values in $PS2$, it would be equivalent to simply push $w_{i}$ to the bottom of $PS2$ if it matched the larger value in the $(2,1)$-pop stack of depth $2$ rather than keep track of how many smaller entries there are. However, we present the algorithm in this form for the option of generalization if desired.
Algorithm 3 $(2,1)$-Pop Stack of Depth $2$ Sorting Algorithm
$w=w_{1}w_{2}\cdots w_{n}$ is a word of length $[n]$.
$PS2$ is a $(2,1)$-Pop Stack of Depth $2$
$O$ is an empty array.
$i=1$ is the index of the next entry in the input
$j=0$ is the counter for the number of copies of the smallest value in the $(2,1)$-Pop Stack of Depth $2$
procedure PopStackSort2($w,n$)
     while $i<n+1$ do
         if $PS2$ is empty then
              push $w_{i}$ to the top of $PS2$
              $i=i+1$
              $j=1$
         else if $w_{i}=$ the top/smallest element(s) in $PS2$ then
              push $w_{i}$ to the top of $PS2$
              $i=i+1$
              $j=j+1$
         else if $w_{i}=$ the larger element(s) in $PS2$ then
              push $w_{i}$ to the $(j+1)$st position of $PS2$
              $i=i+1$
         else if there are already two distinct values in $PS2$ then
              while $PS2$ is nonempty do
                  pop the top entry of $PS2$ to the end of $O$
              end while
         else if $p_{i}<$ top element of $PS2$ then
              push $p_{i}$ to the top of $PS2$
              $i=i+1$
              $j=1$
         else
              while $PS2$ is nonempty do
                  pop the top entry of $PS2$ to the end of $O$
              end while
         end if
     end while
return $O$
end procedure
Theorem 3.23.
Algorithm 3 is optimal for sorting words with an $(r,1)$-pop stack of depth $2$.
Proof.
By way of contradiction let $w$ be a sortable word that is not sorted by our algorithm on a $(2,1)$-pop stack of depth 2. Specifically, $S(w)$ must contain a descent, say $w_{j}w_{k}$. Because Algorithm 3 never allows a descent in the pop stack, $w_{j}$ must have been the last entry popped from one pop operation and $w_{k}$ must have been the first entry popped from the next pop operation. The entries $w_{j}$ and $w_{k}$ must have appeared in the same relative order (forming an inversion) in $w$. The only way to remove this inversion would be for both entries to be in the pop stack at the same time.
Notice Algorithm 3 only pops the pop stack in two cases. In the first case, the pop stack already contained two distinct values that are different from the next input value. In this case, $w_{j}$ is one of the larger entries in the pop stack. Also, $w_{j}$ is the first entry in $w$ from the all of entries of the same value popped at the same time as $w_{j}$ because a new entry only enters in the pop stack if it is smaller than the other value represented or the new entry matches a value already there and is placed above it in the pop stack. In fact, $w_{j}$ must be pushed into an empty pop stack as we are not inserting any unmatched larger entries into the pop stack after a smaller entry. Thus there must be a smaller entry $w_{m}$ popped at the same time as $w_{j}$ and was inserted into the pop stack after $w_{j}$. Notice $w_{j}w_{m}w_{k}$ form a $201$ or $210$ pattern in $w$ which is forbidden.
Otherwise, when $w_{j}$ is popped from the pop stack, only values matching $w_{j}$ can be in the pop stack. For the next entry $w_{i}$, not to enter the pop stack, we must have $w_{j}<w_{i}$. In this case, $w_{j}w_{i}w_{k}$ form a $120$ pattern in $w$ which is also forbidden. ∎
The enumeration problem for this $(2,1)$-pop stack of depth $2$ can be considered for permutations, words, and specifically inversion sequences. The permutations of length $n$ avoiding $231,312,321$, i.e. $120,201,210$, are known as the layered permutations where each layer can have size one or two. Note that layered permutations are exactly the permutations that avoid $231,312$. Further, the restriction of avoiding $321$ forces each layer (consecutive decreasing subsequence) to have size at most two. These restricted layered permutations are enumerated by the Fibonacci numbers as shown by Simion and Schmidt [26].
Corollary 3.24.
[…]
Burstein [5, Theorem 5.1], enumerated the words of length $n$ over alphabet $[k]$ avoiding $120,201,210$.
Theorem 3.25.
(Burstein) The generating function for the number of words of length $n$ over alphabet $[k]$ sortable by a $(2,1)$-pop stack of depth $2$ is given by the coefficient of $x^{n}y^{k}$ in $F(x,y)$ where
$F(x,y)=\frac{(1-x)(1-2x)-((1-x)(1-2x)+x^{2})y}{(1-x)(1-2x)-(1-x)(2-3x)y+(1-2x)y^{2}}.$
The case for the enumeration of sortable inversion sequences by a $(2,1)$-pop stack of depth $2$ (in the form of avoidance of $120,201,210$ or otherwise) does not appear in the literature prior to this work, but can be solved using generating trees as demonstrated in subsection 4.1. Specifically, the generating function for these sortable inversion sequences is given in Theorem 4.3.
4.   The generating tree method
4.1 A formula for the generating function for the number of inversion sequences of length $n$ that avoid $120,201,210$
To enumerate the inversion sequences of length $n$ avoiding $120,201,210$, i.e. the inversion sequences sortable by a $(2,1)$-pop stack of depth $2$, we use the generating tree method shown by Kotsireas, Yıldırım, and the first author [16].
Let ${\mathbf{A}}=\cup_{n=0}^{\infty}{\mathbf{I}}_{n}(120,201,210)$ be the set of all nonempty inversion sequences that avoid $120$, $201$, and $210$. We construct a pattern-avoidance tree ${\mathbf{T}}$ for the class of pattern-avoiding inversion sequences ${\mathbf{A}}$ as follows.
  1. 1.
    The root is 00 (inversion sequence with one letter), that is, $0\in{\mathbf{T}}$ at level $1$.
  2. 2.
    Recursively construct the nodes at level $n+1$ of tree ${\mathbf{T}}$ from the nodes at level $n$ by inserting a new letter at the end of the inversion sequence. That is, the children of $e=e_{0}\cdots e_{n}\in{\mathbf{I}}_{n}\cap{\mathbf{A}}$ are the inversion sequences $e^{*}=e_{0}\cdots e_{n}j$ with $j=0,1,\ldots,n+1$ where $e^{*}\in{\mathbf{I}}_{n+1}\cap{\mathbf{A}}$.
Now, we relabel the vertices of the tree ${\mathbf{T}}$ as follows. Define ${\mathbf{T}}(e)$ to be the subtree consisting of the inversion sequence $e$ as the root and its descendants in ${\mathbf{T}}$. We say that $e$ is equivalent to $e^{\prime}$, denoted by $e\sim e^{\prime}$, if and only if ${\mathbf{T}}(e)\cong{\mathbf{T}}(e^{\prime})$ (in the sense of plane trees). Let ${\mathbf{T}}^{\prime}$ be the same tree ${\mathbf{T}}$ where we replace each node $e$ by the first node $e^{\prime}\in{\mathbf{T}}$ from top to bottom and from left to right in ${\mathbf{T}}$ such that ${\mathbf{T}}(e)\cong{\mathbf{T}}(e^{\prime})$.
Definition 4.1.
A succession rule, denoted $x\rightsquigarrow y_{1},y_{2},\ldots,y_{k}$, will indicate that the set $\{y_{1},y_{2},\ldots,y_{k}\}$ will be such that the generating subtrees rooted at $y_{1},y_{2},\ldots,y_{k}$ are equivalent to those rooted at $z_{1},z_{2},\ldots,z_{k}$ where $\{z_{1},z_{2},\ldots,z_{k}\}$ is the complete set of children of $x$.
Lemma 4.2.
The generating tree ${\mathbf{T}}^{\prime}$ is given by root 00 and the following succession rules
$a_{m}$ $\rightsquigarrow a_{m+1},b_{m,1},\ldots,b_{m,m},$
$b_{m,j}$ $\rightsquigarrow(b_{m+2-j,1})^{j},b_{m+1,j},b_{m+1-j,1},\ldots,b_{m+1-j,m+1-j},$
where $a_{m}=0^{m}$ and $b_{m,j}=a_{m}j$ for $1\leq j\leq m$.
Proof.
We label the inversion sequence $0\in{\mathbf{I}}_{0}$ by $a_{1}$. Thus, $a_{1}\rightsquigarrow a_{2},b_{1,1}$. More generally, by the definitions, the children of $a_{m}\in{\mathbf{T}}^{\prime}$ are $a_{m}0,a_{m}1,\ldots,a_{m}m$, which can also be denoted $a_{m+1},b_{m,1},\ldots,b_{m,m}$, respectively. Thus, the rule $a_{m}\rightsquigarrow a_{m+1},b_{m,1},\ldots,b_{m,m}$ holds.
Also, the children of $b_{m,j}\in{\mathbf{T}}^{\prime}$ are $b_{m,j}0,b_{m,j}1,\ldots,b_{m,j}(m+1)$. Notice ${\mathbf{T}}(b_{m,j}k)\cong{\mathbf{T}}(b_{m+2-j,1})$ with $k=0,1,\ldots,j-1$. Clearly, when $j=1$, any inversion sequence $0^{m}10\pi^{\prime}$ in ${\mathbf{I}}_{n}(120,201,210)$ if any only if $0^{m+1}1\pi^{\prime}$ in ${\mathbf{I}}_{n}(120,201,210)$. Thus, we can assume that $11\leq j\leq k-$ 1. To prove the claim for such $j$, let $\pi=0^{m}jk\pi^{\prime}\in{\mathbf{I}}_{n}(120,201,210)$. Since $j>k$ and $\pi$ avoids both $201$ and $210$, we see that there are no letters in $\pi^{\prime}$ belonging to the set $\{0,\ldots,k-1,k+1,\ldots,j-1\}$. Thus, by mapping $\pi$ to $0^{m+2-k}1\pi^{\prime\prime}$, where $\pi^{\prime\prime}$ is obtained from $\pi^{\prime}$ by mapping each letter $x$ of $\pi^{\prime}$ to $x+1-j$ with $x\neq k$, and mapping the letter $k$ to 00. Hence, we see that $\pi\in{\mathbf{I}}_{n}(120,201,210)$ if and only if $0^{m+2-j}1\pi^{\prime\prime}\in I_{n+2-j}(120,201,210)$.
As none of our patterns have repeated letters, it is not hard to see ${\mathbf{T}}(b_{m,j}j)\cong{\mathbf{T}}(b_{m+1,j})$. Here we map any inversion sequence $0^{m}jj\pi^{\prime}\in{\mathbf{I}}_{n}$ to $0^{m+1}j\pi^{\prime}\in I_{n}$ with $1\leq j\leq m$, so this map respects the rule of avoiding $120,201,210$.
In the last cases when $k=j+1,j+2,\ldots,m+1$, we have a similar mapping as with the smaller $k$ values. Specifically, ${\mathbf{T}}(b_{m,j}k)\cong{\mathbf{T}}(b_{m+1-j,k-j})$. To see this, again, let $\pi=0^{m}jk\pi^{\prime}\in I_{n}(120,201,210)$. Since $\pi$ avoids $120$, we see that each letter of $\pi^{\prime}$ is at least $j$. Let $\pi^{\prime\prime}$ be the obtained sequence from $\pi^{\prime}$ by decreasing each nonzero letter of $\pi^{\prime}$ by $j$. Then the map from $\pi=0^{m}jk\pi^{\prime}\in{\mathbf{I}}_{n}(120,201,210)$ to $0^{m+1-j}(k-j)\pi^{\prime\prime}\in{\mathbf{I}}_{n-j}(120,201,210)$ is a bijection. Hence, we have the following rule $b_{m,j}\rightsquigarrow(b_{m+2-j,1})^{j},b_{m+1,j},b_{m+1-j,1},\ldots,b_{m+1-j,m+1-j}$, which completes the proof. ∎
Define $A_{m}(x)$ (respectively, $B_{m,j}(x)$) to be the generating function for the number of nodes at level $n\geq 1$ for the subtree of ${\mathbf{T}}(B;a_{m})$ (respectively, ${\mathbf{T}}(B;b_{m,j})$), where the root stays at level 00. Thus, by Lemma 4.2, we have
$A_{m}(x)$ $=x+xA_{m+1}(x)+x\sum_{j=1}^{m}B_{m,j}(x),$ (4)
$B_{m,j}(x)$ $=x+jxB_{m+2-j,1}+xB_{m+1,j}(x)+x\sum_{k=1}^{m+1-j}B_{m+1-j,k}(x),$ (5)
for all $1\leq j\leq m$.
In order to solve the above recurrence relations, we define
$A(v)$ $=\sum_{m\geq 1}A_{m}(x)v^{m-1},$
$B(v,u)$ $=\sum_{m\geq 1}\sum_{j=1}^{m}B_{m,j}(x)u^{m-j}v^{m-1},\text{ and}$
$C(v)$ $=\sum_{m\geq 1}B_{m,1}(x)v^{m-1}.$
Thus Equations (4)-(5) can be written as
$A(v)$ $=\frac{x}{1-v}+\frac{x}{v}(A(v)-A(0))+xB(v,1),$ (6)
$B(v,u)$ $=\frac{x}{(1-v)(1-uv)}+\frac{x}{uv(1-v)^{2}}(C(vu)-C(0))+\frac{x}{uv}(B(v,u)-B(v,0))+\frac{x}{1-v}B(uv,1),$ (7)
$C(v)$ $=\frac{x}{1-v}+\frac{2x}{v}(C(v)-C(0))+xB(v,1),$ (8)
where the last equation (8) is the translating of (5) with $j=1$.
To solve the system from Equations (7)-(8), we make the following guess based on the first terms of the generating functions $B(v,1)$, $C(v)$, and $A(v)$:
$B(v,1)$ $=\frac{1}{(1-v)^{2}}C(v)-\frac{v}{(1-v)^{2}}A(v).$ (9)
Next, we solve the system from Equations (7)-(9), which satisfies the original system from Equations (7)-(8). By substituting the expression of $B(v,1)$ from Equation (9) into Equation (7), and solving for $C(v)$, we obtain
$C(v)$ $=\frac{v^{3}-2v^{2}+2vx+v-x}{vx}A(v)+\frac{(1-v)^{2}}{v}A(0)+v-1.$ (10)
From here, use Equation (9) and Equation (10) to rewrite Equation (8) as
$\frac{2x^{2}-3x(x+1)v+(5x+1)v^{2}-2(x+1)v^{3}+v^{4}}{v^{2}x}A(v)$
$=\frac{2x-(3x+1)v+2(x+1)v^{2}-v^{3}}{v^{2}}A(0)-\frac{2x}{v}C(0)-\frac{(v-2x)(v-1)}{v}.$ (11)
Let $K(v)=2x^{2}-3x(x+1)v+(5x+1)v^{2}-2(x+1)v^{3}+v^{4}$ be the kernel of this equation. Note that for the kernel equation $K(v)=0$ there are four roots, say $v_{1},v_{2},v_{3},v_{4}$, where
$v_{1}$ $=1+\frac{-1+\sqrt{5}}{2}x+\frac{-5+4\sqrt{5}}{5}x^{2}+\cdots,$
$v_{2}$ $=1+\frac{-1-\sqrt{5}}{2}x+\frac{-5-4\sqrt{5}}{5}x^{2}+\cdots,$
$v_{3}$ $=2x+2x^{2}+10x^{3}+\cdots,$
$v_{4}$ $=x-x^{3}-x^{4}+\cdots.$
By taking Equation (11) with either $v=v_{3}$ or $v=v_{4}$, we obtain a system of equations in $A(0)$ and $C(0)$. We solve this system, obtaining
$A(0)$ $=-\frac{(v_{1}+v_{2}-2x-1)v_{1}v_{2}}{v_{1}v_{2}(v_{1}+v_{2})-2(1+x)v_{1}v_{2}+2x},$
$C(0)$ $=-\frac{v_{1}v_{2}(v_{2}-1)(v_{1}-1)-(v_{1}^{2}v_{2}+v_{1}v_{2}^{2}-2v_{1}^{2}-3v_{1}v_{2}-2v_{2}^{2}+2v_{1}+2v_{2})x+2(v_{1}v_{2}-2v_{1}-2v_{2}+2)x^{2}}{2x(v_{1}v_{2}(v_{1}+v_{2})-2(1+x)v_{1}v_{2}+2x)}.$
By using expressions of $A(0),C(0)$, we obtain an explicit formula for $A(v)$ from Equation (11). We can then obtain an explicit formula for $C(v)$ from Equation (10). Then we can use that formula to get an explicit formula for $B(v,1)$ from Equation (9). Finally, we have the information we need to obtain an explicit formula for $B(v,u)$ from Equation (7). We omit the presentations of the expressions $A(v),C(v),B(v,1),B(v,u)$ because they are very lengthy. However, these expressions are satisfy the system of Equations (6)-(8). Hence, we can state the following result.
Theorem 4.3.
The generating function for the number of inversion sequences in ${\mathbf{I}}_{n}$ that avoid $120,201,210$ and thus are sortable by a $(2,1)$-pop stack of depth $2$ is given by
$A(0)$ $=-\frac{(v_{1}+v_{2}-2x-1)v_{1}v_{2}}{v_{1}v_{2}(v_{1}+v_{2})-2(1+x)v_{1}v_{2}+2x}$
$=x+2x^{2}+6x^{3}+23x^{4}+101x^{5}+484x^{6}+2468x^{7}+13166x^{8}+72630x^{9}+411076x^{10}+2374188x^{11}$
$+13938018x^{12}+82932254x^{13}+499031324x^{14}+3031610924x^{15}+18568429963x^{16}+114541486785x^{17}$
$+710973143614x^{18}+4437415155234x^{19}+27831038618735x^{20}+175318861863701x^{21}$
[…]
4.2 The generating tree for the pop stack sortable inversion sequences
As in other cases, one can show that the generating tree (based on Algorithm [16]) of $I_{n}(120,201,1010)$, that is, of the pop stack sortable inversion sequences, is given by a root $a_{1}$ and the rules
$a_{m}$ $\rightsquigarrow a_{m+1},b_{m,1},\ldots,b_{m,m},$
$b_{m,j}$ $\rightsquigarrow c_{m+1-j,1},\ldots,c_{m,j},b_{m+1,j},b_{m+1-j,1},\ldots,b_{m+1-j,m+1-j},$
$c_{m,j}$ $\rightsquigarrow c_{m+2-j,1},\ldots,c_{m+1,j},a_{m+3-j},b_{m+2-j,1},\ldots,b_{m+2-j,m+2-j},$
where $a_{m}=0^{m}$, $b_{m,j}=0^{m}j$ and $c_{m,j}=0^{m}j(j-1)$.
For any sequence of nodes $f_{m,j}$, we define $F_{m,j}(x)$ as the generating functions for the number of nodes in the subtrees $\mathcal{T}(B;f_{m,j})$, where $f\in\{a,b,c,d\},F\in\{A,B,C\}$. Define
$A(v)$ $=\sum_{m\geq 1}A_{m}(x)v^{m-1},$
$B(v,u)$ $=\sum_{m\geq 1}\sum_{i=1}^{m}B_{m,i}v^{m-1}u^{m-i},$
$C(v,u)$ $=\sum_{m\geq 1}\sum_{i=1}^{m}C_{m,i}v^{m-1}u^{m-i}.$
Then the generating function for the case 120,201,1010 is $\frac{x}{1-x}+xB(x,1)$, where $B(x,v)$ satisfies
$A(v)$ $=\frac{x}{1-v}+\frac{x}{v}(A(v)-A(0))+xB(v,1),$
$B(v,u)$ $=\frac{x}{(1-v)(1-uv)}+\frac{x}{uv}(B(v,u)-B(v,0))+\frac{x}{1-v}C(v,u)+\frac{x}{1-v}B(vu,1),$
$C(v,u)$ $=\frac{x}{(1-v)(1-uv)}+\frac{x}{vu(1-v)}(C(v,u)-C(v,0))$
$+\frac{x}{v^{2}u^{2}(1-v)}(A(uv)-A(0)-\frac{uv}{x}A(0)+vu+vuB(vu,1)).$
By applying this system $20$ times starting from $A(x,v)=B(x,v,u)=C(x,v,u)=0$, we obtain the first $20$ coefficients of $A(0)$ as
$A(0)$ $=x+2x^{2}+6x^{3}+23x^{4}+101x^{5}+485x^{6}+2488x^{7}+13414x^{8}+75126x^{9}+433546x^{10}$
$+2563335x^{11}+15461646x^{12}+94835817x^{13}+589997530x^{14}+3715451178x^{15}$
$+23645541066x^{16}+151874732111x^{17}+983428159871x^{18}$
$+6413887925931x^{19}+42100271440339x^{20}+\cdots.$
5.   Open Problems
There are many directions left to continue the study of sortable inversion sequences. The most obvious question is the elusive enumeration problem for the inversion sequences of length $n$ which avoid $120$. Already interesting as a standalone pattern avoidance question, the fact that these inversion sequences are exactly those sortable by the classic stack sorting algorithm by Knuth [15] increases the intrigue.
Open Question 5.1.
How many inversion sequences of length $n$ are stack sortable, i.e. avoid $120$?
We have classified the pop stack sortable inversion sequences and words and provided a generating tree for the inversion sequence case in subsection 4.2. However, the combinatorial enumeration is still open.
Open Question 5.2.
How many inversion sequences of length $n$ are pop stack sortable, i.e. avoid $120,201,1010$?
And the cases restricting pop stacks to particular depths in an analogous way to  [10, 11, 12, 13] and/or the expansion in allowable pushes analogous to  [2] can be interesting for words and inversion sequences.
Open Question 5.3.
Classify and enumerate the inversion sequences of length $n$ that are $(r,1)$-pop stack sortable for pop stacks of depth $k$ for $r\geq 2$ and/or $k\geq 2$.
Another avenue of exploration is considering other sorting machines or algorithms. For instance, the authors considered a sorting algorithm on a single stack prioritizing outputting the next correct entry of the sequence until no other moves were available and then either returning the elements in order to be sorted again [22] or reversing the output for the next pass [23]. In particular, the $2$-reverse-pass sortable permutation class has a reasonably small basis, namely $\{2413,2431,23154\}=\{1302,1320,12043\}$ that may prove tractable.
Open Question 5.4.
How many inversion sequences of length $n$ are $2$-reverse-pass sortable?